# Discrete Fourier Transform versus Discrete Chi-Square Method

### v1

@@ 4.2 Uniqueness ( $\mathrm{C}_{2}$)
We make the most of the Gauss-Markov theorem. When the numerical values of the tested frequencies $\bm{\beta}_{I}$ (Equation 9) are fixed, the model becomes linear and the solution for the other free parameters $\bm{\beta}_{II}$ is unique (Equation 10). All possible $\bm{\beta}_{I}$ frequency combinations are tested (Equation 11). For every tested frequency combination $\bm{\beta}_{I}$, the linear model gives a unique value for the test statistic $z$ (Equations  15 and 16). From all tested frequency combinations, we select the best frequency combination $\bm{\beta}_{\mathrm{I,best}}$ which minimises $z$. The linear model for this best frequency combination $\bm{\beta}_{\mathrm{I,best}}$ gives unique values for the remaining other free parameters $\bm{\beta}_{\mathrm{II,best}}$. The only goal for the massive DCM search (Equation 45) is to find these unique initial free parameter values $\bm{\beta}_{\mathrm{initial}}=[\bm{\beta}_{\mathrm{I,best}},\bm{\beta}_{\mathrm{II,best}}]$ for the non-linear iteration that gives the unique final free parameter values $\bm{\beta}_{\mathrm{final}}$ (Equation 21).
@@ 4.3 Stability ( $\mathrm{C}_{3}$)
We conclude that our numerical DCM model solution fulfills the $\mathrm{C}_{1}$, $\mathrm{C}_{2}$ and $\mathrm{C}_{3}$ conditions of the solution for a well-posed problem (Sections 4.1 - 4.3).
@@ 4.4 Performance
The above-mentioned Model 2, 5 and 7 simulations also reveal that the relative accuracy of amplitude and period estimates is lower than the relative accuracy of signal minimum and maximum epoch estimates (Tables 2, 5 and 7). This statistical effect is the same when DCM is applied to the data of any arbitrary phenomenon. This effect would, for example, explain why our solar cycle amplitude predictions are less accurate than our solar cycle minimum and maximum epoch predictions (Jetsu 2025). DCM detects the correct simulated period values, but DFT detects less accurate period values, which are not always correct. This would explain the inexact signal periods that DFT has detected earlier from the sunspot data [e.g., 34].
We show that the DCM signal detections depend only on $n$ and $\mathrm{SN}$}, not on the time span $\Delta T$. This dimensional effect makes sense. If the data are inside an area $\Delta T\times\Delta y$, the ratio $y_{i}/\sigma_{i}$ reveals or hides the $g_{i}$ model details. DCM is an ideal forecasting method because the $\Delta T$ time span of data is irrelevant. This may motivate fast sampling of large high quality datasets and the use of numerical DCM parallel Python compution code in the analysis of those datasets.

### v2

@@ 4.2 Uniqueness ( $\mathrm{C}_{2}$)
We make the most of the Gauß-Markov theorem. When the numerical values of the tested frequencies $\bm{\beta}_{I}$ (Equation 9) are fixed, the model becomes linear and the solution for the other free parameters $\bm{\beta}_{II}$ is unique (Equation 2.1). All possible $\bm{\beta}_{I}$ frequency combinations are tested (Equation 11). For every tested frequency combination $\bm{\beta}_{I}$, the linear model gives a unique value for the test statistic $z$ (Equations  15 and 16). From all tested frequency combinations, we select the best frequency combination $\bm{\beta}_{\mathrm{I,best}}$ which minimises $z$. The linear model for this best frequency combination $\bm{\beta}_{\mathrm{I,best}}$ gives unique values for the remaining other free parameters $\bm{\beta}_{\mathrm{II,best}}$. The only goal for the massive DCM search (Equation 45) is to find these unique initial free parameter values $\bm{\beta}_{\mathrm{initial}}=[\bm{\beta}_{\mathrm{I,best}},\bm{\beta}_{\mathrm{II,best}}]$ for the non-linear iteration that gives the unique final free parameter values $\bm{\beta}_{\mathrm{final}}$ (Equation 21).
@@ 4.3 Stability ( $\mathrm{C}_{3}$)
We conclude that our numerical DCM model solution fulfils the $\mathrm{C}_{1}$, $\mathrm{C}_{2}$ and $\mathrm{C}_{3}$ conditions of the solution for a well-posed problem (Sections 4.1 - 4.3).
@@ 4.4 Performance
The above-mentioned Model 2, 5 and 7 simulations also reveal that the relative accuracy of amplitude and period estimates is lower than the relative accuracy of signal minimum and maximum epoch estimates (Tables 2, 5 and 7). This statistical effect is the same when DCM is applied to the data of any arbitrary phenomenon. This effect would, for example, explain why our solar cycle amplitude predictions are less accurate than our solar cycle minimum and maximum epoch predictions (Jetsu 2025). DCM detects the correct simulated period values, but DFT detects less accurate period values, which are not always correct. The leakage of DFT spectral power [20, 13] would explain why correct signal periods have not been detected earlier from the sunspot data.
We show that the DCM signal detections depend only on $n$ and $\mathrm{SN}$}, not on the time span $\Delta T$. This dimensional effect makes sense. If the data are inside an area $\Delta T\times\Delta y$, the ratio $y_{i}/\sigma_{i}$ reveals or hides the $g_{i}$ model details. DCM is an ideal forecasting method because the $\Delta T$ time span of data is irrelevant. This may motivate fast sampling of large high quality datasets and the use of numerical DCM parallel Python computation code in the analysis of those datasets.
4.5 Methodological contribution
What is our methodological contribution? DCM has already been published and DFT has been applied in innumerable studies. We would get the same results if DCM were compared to any other time series analysis method than DFT.
Practical time series analysis applications, like digital signal processing, require fast algorithms [20]. Our statistical numerical solution for the non-linear time series analysis model requires a lot of CPU. Yet, the CPU is beside the point if a well-posed numerical solution can be found. DCM proceeds through two stages. For every chosen alternative model, DCM performs a massive number of linear least squares fits (Equation 18) to find the unique initial free parameter values for the non-linear iteration (Equation 21). This first stage devours CPU. The second fast stage, the Fisher-test, then reveals the best model alternative (Equation 39).
DCM is an extremely simple time series analysis method that works in the time domain and uses real numbers. The more elaborate spectral estimating techniques, like DFT, work in the frequency domain and use complex numbers.
We also want to draw attention to our idea of frequency space symmetry. This symmetry reduces the number of tested frequency combinations (Equation 11). For example, were this symmetry not exploited, the four signal $K_{1}=4$ search would give $4!=24$ same solutions in four-dimensional frequency space. The identification of the best frequency combination would be a daunting exercise, let alone the CPU consumed in testing all possible frequency combinations inside a tesseract (a four-dimensional cube). Our one-dimensional periodogram slices can "see inside” such multi-dimensional structures (Equation 20).
Generally, different time series analysis methods suffer from different limitations. We list some of those limitations below.
  1. 1.
    The required data spacing is even or uneven.
  2. 2.
    The number of signals is unknown.
  3. 3.
    The presence and shape of trend are unknown.
  4. 4.
    The degree of noise is unknown.
  5. 5.
    The sample window causes leakage.
  6. 6.
    The leakage weakens frequency resolution.
  7. 7.
    The signal frequencies are too close.
  8. 8.
    The signal periods exceed the data time span.
  9. 9.
    The signal shapes are not pure sinusoids.
  10. 10.
    The forecasts of non-linear models fail.
DCM does not suffer from any of these limitations. This is our methodological contribution.
We hope that our results are considered to represent advances in the field of time series analysis. Our DCM can not fail if the Gauß-Markov theorem is valid.